# Constraining the modified friction in gravitational wave propagation with precessing black hole binaries

### v2

@@ Abstract
A broad class of modified gravities can result in a modified friction effect in the propagation of gravitational waves (GWs). This effect changes the amplitude-damping rate of GWs during their propagation in the cosmological distance and thus modifies the standard luminosity distance of GWs in general relativity. Therefore, one can constrain this modified friction by measuring both the luminosity distance and redshift of the GW sources. In this paper, we investigate the prospects of constraining such modified friction effect by using the precessing binary black holes with ground-based GW detectors. For this purpose, we consider 20 precessing events detected by the GW detector network consisting of two LIGO detectors and two third-generation GW detectors (the Einstein Telescope and the Cosmic Explorer). The redshift information of these events is obtained by identifying their possible host galaxies in the GLADE+ galaxy catalog. We show that the precession in the binary system can improve significantly the precision of luminosity distance and thus lead to a tighter constraint on the modified friction. By assuming narrow priors on cosmological parameters that are consistent with the uncertainties of Planck 2018 results, our analysis shows that the modified friction effect, characterized by two parameters $(\Xi_{0},n)$, can be constrained to be $\Xi_{0}=1.002^{+0.004}_{-0.004}$ and $n=3.257^{+2.595}_{-2.192}$, in which the result of $\Xi_{0}$ is about two orders of magnitude better than current result from an analysis with GWTC-3. Our result sets the stage for future research with third-generation GW detectors, offering new insights into gravitational parameter modifications. It also contributes to the understanding of the properties and applications of binary black hole systems with precession.
@@ I Introduction
The field of gravitational wave (GW) astronomy has undergone remarkable growth since the landmark detection of GW150914, a GW event stemming from the merger of two massive black holes, by the LIGO-Virgo collaboration in 2015 gw150914 . This groundbreaking discovery opens a new era in gravitational physics, and later, the LIGO-Virgo-KAGRA (LVK) scientific collaboration has identified about 90 GW events in the subsequent years gw-other ; gw170817 ; LIGOScientific:2017ycc ; gwtc1 ; gwtc2 ; LIGOScientific:2021djp . These observations are consistent with the gravitational waves generated by the mergers of compact binary systems, as predicted by general relativity (GR). To this end, gravitational waves have proven to be an indispensable tool for testing GR under the extreme conditions of strong gravitational fields and dynamical environments, thereby cementing their significance in the era of GW astronomy gw150914-testGR ; gw170817-testGR ; gw170817-speed ; testGR_GWTC1 ; testGR_GWTC2 . In stride with the advancements made by the LVK, the next generation of ground-based GW detectors including the Einstein Telescope (ET), the Cosmic Explorer (CE) intro1 ; intro2 and space-based ones including LISA Robson:2018ifk ; LISACosmologyWorkingGroup:2022jok , Taiji Ruan:2018tsw ; Hu:2017mde , and Tianqin Liu:2020eko ; Milyukov:2020kyg , are currently in the planning stages. These forthcoming detectors are expected to play a pivotal role in further propelling the frontier of gravitational physics.
@@
While GR stands as the most successful theory of gravity, it encounters substantial difficulties in explaining certain phenomena, including theoretical obstacles like singularities and the challenge of quantizing gravity, as well as observational enigmas such as dark matter and dark energy in7 ; in66 . To confront these issues, a comprehensive array of experiments has been carefully crafted to rigorously test GR’s theoretical predictions. However, these experiments primarily scrutinize the theory’s implications in the weak-field limit in3 ; in4 ; in5 . GWs, on the other hand, arise from regions of strong and highly dynamical gravitational fields and interact weakly with matter, positioning them as powerful tools for probing the nature of gravity in6 . In light of the GW signals we have detected, a spectrum of alternative gravitational theories has been proposed to address enduring queries within the GR paradigm MG1 ; MG2 ; MG3 . Consequently, the tests of the modified gravities are crucial to ascertain the definitive theory of gravitational interaction.
@@
A broad class of modified gravities models can introduce a modified friction effect in the propagation of GWs in17 ; in18 ; in19 ; in8 ; in9 ; in10 ; in11 ; in12 ; in13 ; in14 ; in15 ; in16 , see also Table. 1 for a list of modified theories with such effect. This effect, when it is frequency-independent, affects the amplitude-damping rate of GWs during their propagation in the cosmological distance, which in turn changes the standard luminosity distance of GWs in GR. Thus, it is possible to test such an effect with the standard sirens, by measuring both the luminosity distance and the redshift of GW sources. Constraints on the modified friction effect have been derived through multi-messenger analysis of GW170817, the only binary neutron star merger detected by the LVK Collaboration to date, for which the redshift has been directly measured from the corresponding electromagnetic counterparts work1 ; work4 ; Mastrogiovanni:2020mvm . Several other constraints have also been obtained by using GW events with redshifts information inferred from the host galaxies Finke:2021aom ; Mancarella:2022cgn ; Chen:2023wpj , or BBH mass function Mancarella:2021ecn ; Leyde:2022orh ; Ezquiaga:2021ayr ; Mancarella:2022cgn . Recently, several forecasts for constraining the modified GW friction with future GW detectors have also been carried out in detail, see refs. Liu:2023onj ; work10 ; Chen:2024xkv ; Narola:2023viz ; Matos:2022uew and references therein.
@@
To constrain the modified GW friction, it is essential to measure the luminosity distance and the redshift of the GW source accurately and independently. For a GW source without an electromagnetic counterpart, an alternative technique is to estimate the source redshift using the statistical redshift information of the possible host galaxies in the galaxy catalogs. To narrow down the number of galaxies in the volume, one needs to improve measurement precision of both the luminosity distance and the sky location of the source. It is recently shown in ref. Raymond:2008im ; Yun:2023ygz ; Green:2020ptm that the accuracy of luminosity distance estimations can be significantly improved through the observation of GW events from the mergers of precessing binary black holes, and thus can be better candidates for standard sirens Yun:2023ygz . The presence of the procession in the binary black hole system has been observed in several GW events in GWTC-3 Hannam:2021pit ; Islam:2023zzj and can help in breaking the distance-inclination degeneracy to improve the precision of the distance measurement Raymond:2008im ; Yun:2023ygz ; Green:2020ptm . Moreover, a GW network through the integration of multiple interferometers, both operational and anticipated to be functional (CE and ET), can considerably refine the localization precision of GW events. This enables the simultaneous improvement in the measurement precision of both luminosity distances and redshifts. Consequently, by synergizing the aforementioned precession mechanism with a third-generation ensemble of GW detectors, we stand on the brink of obtaining high-quality luminosity distance and redshift data, which promises to advance the constraints on GW friction beyond the achievements of preceding studies.
@@
In this paper, we investigate the prospects of constraining such modified friction effect by using the precessing binary black holes with a ground-based GW detector network consisting of two LIGO detectors and two third-generation GW detectors (ET and CE). We generated GW signals from merging BBH systems exhibiting precession, utilizing the IMRPhenomPv3 waveform model 4 . These signals were then subjected to a simulated detection process incorporating the combined sensitivities of the two LIGO interferometers, CE and ET. We then determine the redshifts of the host galaxies associated with these GW events within the GLADE+ catalog, leveraging the enhanced spatial localization capabilities of third-generation GW detectors. It is shown that the precession in the binary system can improve significantly the precision of luminosity distance and thus lead to a tighter constraint on the modified GW friction.
@@
This paper is organized as follows. We briefly introduce the theory related to modified gravity and the GW friction and give the method of the $c_{M}$-parametrization. In Sec. III, we simulate the GW signal with precession and obtain the luminosity distance from interferometer array measurements. Sec. IV is dedicated to the construction of the error box and the determination of the redshift distribution via host galaxies. Sec. V introduces the fundamental statistical framework utilized in this study, encompassing Bayesian analysis and the MCMC method, followed by the presentation of GW friction parameter estimation results in Sec. VI. The paper concludes with a summary and discussion in Sec. VII.
@@
Throughout this paper, the metric convention is chosen as (,+,+,+)(-,+,+,+)( - , + , + , + ), and greek indices $(\mu,\nu,\cdot\cdot\cdot)$ run over $0,1,2,3$ and latin indices $(i,\;j,\;k)$ run over $1,2,3$. We choose the units to $G=c=1$.
II Modified GW friction in the gravitational propagations
@@
In this section, we present a brief review of the modified frictions of GWs which can modify the damping rates of the two tensorial modes of GWs.
@@ II.1 Modified GW propagations
where $\tau$ denotes the conformal time, which relates to the cosmic time $t$ by $dt=ad\tau$, and $a$ is the scale factor of the universe. Throughout this paper, we set the present scale factor $a_{0}=1$. $h_{ij}$ denotes the GWs, which represent the transverse and traceless metric perturbations, i.e,
@@
To the modified frictions in GW propagations, it is convenient to decompose the GWs into circular polarization modes. To study the evolution of $h_{ij}$, we expand it over spatial Fourier harmonics,
@@
where $e_{ij}^{A}$ denote the circular polarization tensors and satisfy the relation
@@
In such a parametrization, the new effects arising from theories beyond GR are fully characterized by four parameters: $\bar{\nu}$, $\bar{\mu}$, $\nu_{A}$, and $\mu_{A}$. Different parameters correspond to different effects on the propagation of GWs. These effects can be divided into three classes: 1) the frequency-independent effects induced by $\bar{\mu}$ and $\bar{\nu}$ which include the modification to the GW speed and friction; 2) the parity-violating effects induced by $\nu_{A}$ and $\mu_{A}$ which include the amplitude and velocity birefringences of GWs respectively; and 3) the Lorentz-violating effects induced by frequency-dependent $\bar{\nu}$ and $\bar{\mu}$ which include the frequency-independent damping and nonlinear dispersion relation of GWs respectively. The corresponding modified theories with specific forms of the four parameters ${\cal H}\bar{\nu}$, $\bar{\mu}$, ${\cal H}\nu_{A}$, and $\mu_{A}$ are summarized in Table I in ref. Zhu:2023wci . The parity- and Lorentz-violating effects, in general, induce amplitude or phase corrections to the GW waveforms of the compact binary inspiral systems since these effects are frequency-dependent Zhu:2023wci ; waveform . since they are frequency-dependent, they in general induce refs. Zhu:2023wci ; waveform derived their corrections to the GW waveforms of the compact binary inspiral systems. One then can constrain the parity- and Lorentz-violating effects by comparing the modified waveforms with the GW signals, see refs. Zhang:2024rel ; Zhu:2023wci ; Qiao:2019wsh ; Hou:2024xbv ; Gong:2023ffb ; Zhu:2022uoq ; Niu:2022yhr ; Gong:2021jgg ; Wu:2021ndf ; Wang:2020cub ; Zhao:2019szi and references therein. However, these cases are not in the scope of this research, and in this paper, we only focus on the frequency-independent cases.
Table 1: Corresponding parameters ${\cal H}\bar{\nu}$ and $\bar{\mu}$ in specific modified theories of gravity. A comprehensive list of modified gravities with different types of modified GW propagations can be found in Table. I of ref. Zhu:2023wci .
@@ II.2 Modified GW friction and luminosity distance
When the parameters $\bar{\mu}$ and $\bar{\nu}$ are frequency-independent, they can induce two distinct and frequency-independent effects on the propagation of GWs. One is the modification to the speed of GWs due to the nonzero of $\bar{\mu}$, and another effect is the modified friction term of the GWs if $\bar{\nu}$ is nonzero. These frequency-independent effects can arise from a lot of modified gravities, for example, the scalar-tensor theory, extra dimension, Einstein-Æther theory, etc., as summarized in Table 1.
With parameter $\bar{\mu}$, the speed of the GWs are modified in a frequency-independent manner, $c_{\rm gw}\simeq 1+\frac{1}{2}\bar{\mu}$. For a GW event with an electromagnetic counterpart, $c_{\rm gw}$ can be constrained by comparison with the arrival time of the photons. For the binary neutron star merger GW170817 and its associated electromagnetic counterpart GRB170817A LIGOScientific:2017zic , the almost coincident observation of both the electromagnetic wave and GW places an exquisite bound on $\bar{\mu}$, i.e., $-3\times 10^{-15}<\frac{1}{2}\bar{\mu}<7\times 10^{-16}$. Note that here we set the speed of light $c=1$.
The parameter $\bar{\nu}$ induces an additional friction term on the propagation equation of GWs. In a lot of modified gravities, this term is time-dependent and one can write it in terms of the running of the effective Planck mass in the form of Lagos:2019kds
@@
However, such a probe relies sensitively on the time evolution of $\mathcal{H}\bar{\nu}$, which is generally priorly unknown. There are two approaches to parametrize the time evolution of $\mathcal{H}\bar{\nu}$: the $c_{M}$-parametrization Lagos:2019kds which is based on the evolution of the dark energy in the Universe and the $\Xi$-parametrization Belgacem:2018lbp which is theory-based parametrization that can fit a lot of modified gravities.
For $c_{M}$-parametrization, the GW friction is written as Lagos:2019kds
@@
Such parametrization corresponds to
@@
The relation between the $\Xi$-parametrization and $c_{M}$-parametrization has been explored in Mancarella:2021ecn . Several constraints on $(\Xi_{0},n)$ have been obtained using GW events with redshifts information inferred from the corresponding electromagnetic counterparts Mastrogiovanni:2020mvm or host galaxies Ezquiaga:2021ayr , or BBH mass function Mancarella:2021ecn . A recent constraint on $(\Xi_{0},n)$ were from an analysis of GW data in GWTC-3 with BBH mass function, which gives Mancarella:2021ecn
@@
In this paper, we will adopt the $\Xi$-parametrization and explore how future GW observations can improve the current constraints.

### v3

@@ Abstract
A broad class of modified gravities can result in a modified friction effect in the propagation of gravitational waves (GWs). This effect changes the amplitude-damping rate of GWs during their propagation in the cosmological distance and thus modifies the standard luminosity distance of GWs in general relativity. Therefore, one can constrain this modified friction by measuring both the luminosity distance and redshift of the GW sources. In this paper, we investigate the prospects of constraining such modified friction effects by using the precessing binary black holes with ground-based GW detectors. For this purpose, we consider 20 precessing events detected by the GW detector network consisting of two LIGO detectors and two third-generation GW detectors (the Einstein Telescope and the Cosmic Explorer). The redshift information of these events is obtained by identifying their possible host galaxies in the GLADE+ galaxy catalog. We show that the precession in the binary system can significantly improve the precision of the luminosity distance and thus lead to a tighter constraint on the modified friction. By assuming narrow priors on cosmological parameters that are consistent with the uncertainties of Planck 2018 results, our analysis shows that the modified friction effect, characterized by two parameters $(\Xi_{0},n)$, can be constrained to be $\Xi_{0}=1.002^{+0.004}_{-0.004}$ and $n=3.257^{+2.595}_{-2.192}$, in which the result of $\Xi_{0}$ is about 2 orders of magnitude better than current results from an analysis with GWTC-3. Our result sets the stage for future research with third-generation GW detectors, offering new insights into gravitational parameter modifications. It also contributes to the understanding of the properties and applications of binary black hole systems with precession.
@@ I Introduction
The field of gravitational wave (GW) astronomy has undergone remarkable growth since the landmark detection of GW150914, a GW event stemming from the merger of two massive black holes, by the LIGO-Virgo Collaboration in 2015 gw150914 . This groundbreaking discovery opened a new era in gravitational physics. In subsequent years, the LIGO-Virgo-KAGRA (LVK) Collaboration has identified about 90 GW events in the subsequent years gw-other ; gw170817 ; LIGOScientific:2017ycc ; gwtc1 ; gwtc2 ; LIGOScientific:2021djp . These observations are consistent with the gravitational waves generated by the mergers of compact binary systems, as predicted by general relativity (GR). To this end, gravitational waves have proven to be an indispensable tool for testing GR under the extreme conditions of strong gravitational fields and dynamical environments, thereby cementing their significance in the era of GW astronomy gw150914-testGR ; gw170817-testGR ; gw170817-speed ; testGR_GWTC1 ; testGR_GWTC2 . In stride with the advancements made by the LVK, the next generation of ground-based GW detectors, including the Einstein Telescope (ET), the Cosmic Explorer (CE) intro1 ; intro2 , and space-based ones such as LISA Robson:2018ifk ; LISACosmologyWorkingGroup:2022jok , Taiji Ruan:2018tsw ; Hu:2017mde , and Tianqin Liu:2020eko ; Milyukov:2020kyg are currently in the planning stages. These forthcoming detectors are expected to play a pivotal role in further propelling the frontier of gravitational physics.
@@
While GR stands as the most successful theory of gravity, it encounters substantial difficulties in explaining certain phenomena, including theoretical obstacles like singularities and the challenge of quantizing gravity, as well as observational enigmas such as dark matter and dark energy in7 ; in66 . To confront these issues, a comprehensive array of experiments has been carefully crafted to rigorously test GR’s theoretical predictions. However, these experiments primarily scrutinize the theory’s implications in the weak-field limit in3 ; in4 ; in5 . GWs, on the other hand, arise from regions of strong and highly dynamical gravitational fields and interact weakly with matter, positioning them as powerful tools for probing the nature of gravity in6 . In light of the GW signals we have detected, a spectrum of alternative gravitational theories has been proposed to address enduring queries within the GR paradigm MG1 ; MG2 ; MG3 . Consequently, the tests of the modified gravities are crucial to ascertaining the definitive theory of gravitational interaction.
@@
A broad class of modified gravity models can introduce a modified friction effect in the propagation of GWs in17 ; in18 ; in19 ; in8 ; in9 ; in10 ; in11 ; in12 ; in13 ; in14 ; in15 ; in16 (see also Table. 1 for a list of modified theories with such an effect). This effect, when it is frequency independent, affects the amplitude-damping rate of GWs during their propagation in the cosmological distance, which in turn changes the standard luminosity distance of GWs in GR. Thus, it is possible to test such an effect with the standard sirens by measuring both the luminosity distance and the redshift of GW sources. The redshift of a GW source can be directly measured if the source has an electromagnetic counterpart. For GW sources without electromagnetic counterparts, the redshift can be inferred using the mass distribution of binary compact objects Taylor:2012db ; Farr:2019twy ; Ezquiaga:2022zkx ; MaganaHernandez:2024uty or by considering potential host galaxies identified in galaxy catalogs DES:2020nay ; Palmese:2021mjm ; Gair:2022zsa . Constraints on the modified friction effect have been derived through multi-messenger analysis of GW170817, the only binary neutron star merger detected by the LVK Collaboration to date, for which the redshift has been directly measured from the corresponding electromagnetic counterparts work1 ; work4 ; Mastrogiovanni:2020mvm . Several other constraints have also been obtained by using GW events with redshift information inferred from the host galaxies Finke:2021aom ; Mancarella:2022cgn ; Chen:2023wpj , or the BBH mass function Mancarella:2021ecn ; Leyde:2022orh ; Ezquiaga:2021ayr ; Mancarella:2022cgn . Recently, several forecasts for constraining the modified GW friction with future GW detectors have also been carried out in detail (see Refs. Liu:2023onj ; work10 ; Chen:2024xkv ; Narola:2023viz ; Matos:2022uew and references therein). Note that the extra dimension can also induce modified luminosity distance of GWs, and the corresponding constraints were explored in Refs. Pardo:2018ipy ; MaganaHernandez:2021zyc .
@@
To constrain the modified GW friction, it is essential to measure the luminosity distance and the redshift of the GW source accurately and independently. For a GW source without an electromagnetic counterpart, as we mentioned, an alternative technique is to estimate the source redshift using the statistical redshift information of the possible host galaxies in the galaxy catalogs. To narrow down the number of galaxies in the volume, one needs to improve the measurement precision of both the luminosity distance and the sky location of the source. It has recently been shown in Refs. Raymond:2008im ; Yun:2023ygz ; Green:2020ptm that the accuracy of luminosity distance estimations can be significantly improved through the observation of GW events from the mergers of precessing binary black holes, and thus can be better candidates for standard sirens Yun:2023ygz . The presence of the precession in the binary black hole system has been observed in several GW events in GWTC-3 Hannam:2021pit ; Islam:2023zzj and can help break the distance-inclination degeneracy to improve the precision of the distance measurement Raymond:2008im ; Yun:2023ygz ; Green:2020ptm . In addition, inclusion of high-order modes (HOMs) in the parameter estimation of GW signals can also help break the distance-inclination degeneracy (see Refs. London:2017bcn ; Mascioli:2025cnw ; Borhanian:2020vyr and references therein). Moreover, a GW network through the integration of multiple interferometers, both operational and anticipated to be functional (CE and ET), can considerably refine the localization precision of GW events. This enables the simultaneous improvement in the measurement precision of both luminosity distances and redshifts. Consequently, by synergizing the aforementioned precession mechanism with a third-generation ensemble of GW detectors, we stand on the brink of obtaining high-quality luminosity distance and redshift data, which promises to advance the constraints on GW friction beyond the achievements of preceding studies.
@@
In this paper, we investigate the prospects of constraining such modified friction effects by using the precessing binary black holes with a ground-based GW detector network consisting of two LIGO detectors and two third-generation GW detectors (ET and CE). We generate GW signals from merging BBH systems exhibiting precession, utilizing the IMRPhenomPv3 waveform model 4 . These signals are then subjected to a simulated detection process incorporating the combined sensitivities of the two LIGO interferometers, CE and ET. We then determine the redshifts of the host galaxies associated with these GW events within the GLADE+ catalog, leveraging third-generation GW detectors’ enhanced spatial localization capabilities. We show that the precession in the binary system can significantly improve the precision of the luminosity distance and thus lead to a tighter constraint on the modified GW friction.
@@
This paper is organized as follows. We briefly introduce the theory related to modified gravity and the GW friction and give the method of $c_{M}$ parametrization. In Sec. III, we simulate the GW signal with precession and obtain the luminosity distance from interferometer array measurements. In Sec. IV, we illustrate that the third-generation detectors have better luminosity distance and precession estimation accuracies compared to the second-generation detectors by simulating three different detector combinations. Section V is dedicated to the construction of the error box and the determination of the redshift distribution via host galaxies. Sec. VI introduces the fundamental statistical framework utilized in this study, encompassing Bayesian analysis and the Markov chain Monte Carlo (MCMC) method, followed by the presentation of GW friction parameter estimation results in Sec. VII. The paper concludes with a summary and discussion in Sec. VIII.
@@
Throughout this paper, the metric convention is chosen as (,+,+,+)(-,+,+,+)( - , + , + , + ), and Greek indices $(\mu,\nu,\cdot\cdot\cdot)$ run over $0,1,2,3$, and Latin indices $(i,\;j,\;k)$ run over $1,2,3$. We choose the units $G=c=1$.
II Modified GW friction in gravitational propagations
@@
In this section, we present a brief review of the modified friction of GWs which can modify the damping rates of the two tensorial modes of GWs.
@@ II.1 Modified GW propagations
where $\tau$ denotes the conformal time, which relates to the cosmic time $t$ by $dt=ad\tau$, and $a$ is the scale factor of the universe. Throughout this paper, we set the present scale factor $a_{0}=1$, and $h_{ij}$ denotes the GWs, which represent the transverse and traceless metric perturbations, i.e,
@@
To study the modified frictions in GW propagations, it is convenient to decompose the GWs into circular polarization modes. To study the evolution of $h_{ij}$, we expand it over spatial Fourier harmonics,
@@
where $e_{ij}^{A}$ denotes the circular polarization tensors and satisfies the relation
@@
In such a parametrization, the new effects arising from theories beyond GR are fully characterized by four parameters: $\bar{\nu}$, $\bar{\mu}$, $\nu_{A}$, and $\mu_{A}$. Different parameters correspond to different effects on the propagation of GWs. These effects can be divided into three classes: 1) the frequency-independent effects induced by $\bar{\mu}$ and $\bar{\nu}$ which include the modification to the GW speed and friction; 2) the parity-violating effects induced by $\nu_{A}$ and $\mu_{A}$ which include the amplitude and velocity birefringence of GWs respectively; and 3) the Lorentz-violating effects induced by frequency-dependent $\bar{\nu}$ and $\bar{\mu}$ which include the frequency-independent damping and nonlinear dispersion relation of GWs respectively. The corresponding modified theories with specific forms of the four parameters ${\cal H}\bar{\nu}$, $\bar{\mu}$, ${\cal H}\nu_{A}$, and $\mu_{A}$ are summarized in Table I in Ref. Zhu:2023wci . The parity- and Lorentz-violating effects, in general, induce amplitude or phase corrections to the GW waveforms of the compact binary inspiral systems since these effects are frequency-dependent Zhu:2023wci ; waveform . One then can constrain the parity- and Lorentz-violating effects by comparing the modified waveforms with the GW signals (see Refs. Zhang:2024rel ; Zhu:2023wci ; Qiao:2019wsh ; Hou:2024xbv ; Gong:2023ffb ; Zhu:2022uoq ; Niu:2022yhr ; Gong:2021jgg ; Wu:2021ndf ; Wang:2020cub ; Zhao:2019szi ; Li:2024fxy and references therein). However, these cases are not in the scope of this research, and in this paper, we only focus on the frequency-independent cases.
Table 1: Corresponding parameters ${\cal H}\bar{\nu}$ and $\bar{\mu}$ in specific modified theories of gravity. A comprehensive list of modified gravities with different types of modified GW propagations can be found in Table I of Ref. Zhu:2023wci .
@@ II.2 Modified GW friction and luminosity distance
When the parameters $\bar{\mu}$ and $\bar{\nu}$ are frequency independent, they can induce two distinct and frequency-independent effects on the propagation of GWs. One is the modification to the speed of GWs due to the nonzero of $\bar{\mu}$, and another effect is the modified friction term of the GWs if $\bar{\nu}$ is nonzero. These frequency-independent effects can arise from many modified gravities, for example, the scalar-tensor theory, extra dimensions, Einstein-Æther theory, etc. as summarized in Table 1.
With parameter $\bar{\mu}$, the speed of the GWs is modified in a frequency-independent manner, $c_{\rm gw}\simeq 1+\frac{1}{2}\bar{\mu}$. For a GW event with an electromagnetic counterpart, $c_{\rm gw}$ can be constrained by comparison with the arrival time of the photons. For the binary neutron star merger GW170817 and its associated electromagnetic counterpart GRB170817A LIGOScientific:2017zic , the almost coincident observation of both the electromagnetic wave and GW places an exquisite bound on $\bar{\mu}$, i.e., $-3\times 10^{-15}<\frac{1}{2}\bar{\mu}<7\times 10^{-16}$. Note that here we set the speed of light as $c=1$.
The parameter $\bar{\nu}$ induces an additional friction term on the propagation equation of GWs. In many modified gravities, this term is time-dependent and one can write it in terms of the running of the effective Planck mass in the form of Lagos:2019kds
@@
However, such a probe relies sensitively on the time evolution of $\mathcal{H}\bar{\nu}$. There are two approaches to parametrize the time evolution of $\mathcal{H}\bar{\nu}$: the $c_{M}$ parametrization Lagos:2019kds which is based on the evolution of the dark energy in the Universe and the $\Xi$ parametrization Belgacem:2018lbp , which is theory-based parametrization that can fit many modified gravities.
For $c_{M}$ parametrization, the GW friction is written as Lagos:2019kds
@@
Such a parametrization corresponds to
@@
The relation between the $\Xi$ parametrization and $c_{M}$ parametrization has been explored in Mancarella:2021ecn . Several constraints on $(\Xi_{0},n)$ have been obtained using GW events with redshift information inferred from the corresponding electromagnetic counterparts Mastrogiovanni:2020mvm or host galaxies Ezquiaga:2021ayr , or BBH mass function Mancarella:2021ecn . A recent constraint on $(\Xi_{0},n)$ was from an analysis of GW data in GWTC-3 with the BBH mass function, which gives Mancarella:2021ecn
@@
In this paper, we will adopt the $\Xi$ parametrization and explore how future GW observations can improve the current constraints.